If x+y=6, xy=9, then x2+y2 is equal to_____?
Answer by mananth(16946)

(x+y)=6
xy=9
(x+y)^2=36
x^2+2xy +y^2=36
xy =9
x^2+18+y^2=36
x^2+y^2=36-18
x^2+y^2=18
m.ananth@hotmail.ca